Grade 12Physical Chemistry

The V.P of methyl and ethyl alcohol at 25°c are 90 and 40mm of hg respectively. If 12g of CH3OH and 48g of C2H5OH are mixed calculate the V.P exerted by each component the total vapour pressure of mixture and the composition of mixture?

Govind Nair
8 Years agoGrade 12

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Please find the answer to your question
 
Ptotal = pA + pB
Molecular weight of CH3OH = 12 + 3 + 16 + 1 = 32
Molecular weight of C­2H5OH = 24 + 5 + 16 + 1 = 46
According to Raoult’s law
Ptotal = p1 + p2
Where Ptotal = Total vapour pressure of the solution
P1 = Partial vapour pressure of one component
P2 = Partial vapour pressure of other componenet
Again, p1 = Vapour pressure (po1) * mole fraction
Similarly, p2 = Vapour pressure (po2) * mole fraction
Mole fraction of CH3OH = 12/32 /(12/32+48/46) = 0.264
Mole fraction of ethanol = 48/46 /(48/46+12/32) = 0.735
NOTE THIS STEP : Thus now let us first calculate the partial vapour pressures, i.e., p1 and p2 of the two component.
Partial vapour pressure of CH3OH(p1)
= 90 * 0.264 = 23.76 mm
∴ Total vapour pressure of the solution
= 29.4 + 23.76 mm = 53.16 mm
Mole fraction of CH3OH in vapour = 23.76/53.16 = 0.44